# Beyond Grid-Locked Voxels: Neural Response Functions for Continuous Brain Encoding

### v1

@@ 5 Results
Figure 2: Prediction accuracy (Pearson correlation) in low data regime. a. Single-subject models. NRF consistently outperforms baseline models when trained on limited samples from scratch, highlighting the benefit of its continuous mapping. Results are shown for the average median voxel correlation across four subjects, with error bars indicating the standard error of the mean (SEM). b. Cross-subject transfer. Voxel-level prediction accuracy visualized on the cortical surface of subject 7. When pretrained base models from other subjects are available, the NRF finetune ensemble further improves performance over NRF scratch and baselines, showing clear gains across visual regions.
@@ 5.1 Individual subjects encoding
We first evaluated NRF’s neural prediction capability for single-subject data. Training a separate model for each of the 4 subjects and comparing the average neural prediction accuracy across subjects. For comparison, we selected two representative encoding models as baseline comparison. The linear regression model from the BrainDIVE Luo et al. (2023) and the fWRF (Feature-Weighted Receptive Field) encoder St-Yves & Naselaris (2018). Details about the baseline model are in A.3. We also took the result for full data encoding performance from MindSimulator Bao et al. (2025).
We first evaluate NRF under limited-data conditions, since practical applications rarely have access to the tens of thousands of trials collected in large-scale datasets such as NSD. As shown in Figure 2a, NRF achieves significantly higher accuracy than baseline models when trained on small numbers of images. Remarkably, with only 200 training samples, NRF outperforms baselines trained on more than 800 images. We attribute this data efficiency to the anatomical awareness of NRF: by conditioning on spatial coordinates, the model can exploit the smoothness of fMRI responses and learn more effectively from scarce data. This neuroscience-inspired design makes NRF particularly well-suited for realistic, low-data regimes. We then evaluate NRF in the full-data setting, where models were trained on $\sim$9k images. The quantitative evaluation results are shown in Table 1.NRF outperforms baselines on voxelwise prediction metrics and achieves comparable performance on semantic-level evaluations. Figure 3 shows the visualization results of the decoded images. Low-level details and the semantic information of the images are very similar to the ground truth. These results demonstrate that NRF maintains high voxel-level accuracy while also preserving semantic information, confirming its effectiveness across both limited and full data regimes.
In addition, we observed that fWRF achieves unusually high semantic-level scores, in some cases even surpassing reconstructions from measured fMRI. We attribute this to decoder bias: fWRF outputs, while less neurally accurate, may align more closely with the pretrained decoder’s distribution, inflating semantic metrics. These results indicate that semantic-level metrics should be interpreted as a coarse indication of reconstruction quality rather than a strict basis for comparing encoding models.
@@
Table 1: Evaluation results of fMRI prediction accuracy for the model trained on the full dataset. Our NRF model achieves state-of-the-art performance on prediction accuracy and comparable results on semantic metrics. All metrics are calculated across 4 subjects.
@@ 5.2 New Subject adaptation
More importantly, NRF enables cross-subject transfer, allowing knowledge learned from one subject to be adapted to new subjects—a critical property given that collecting fMRI data for new individuals is both resource-intensive and time-consuming. To evaluate this capability, we tested adaptation with 20, 200, and 800 images, corresponding to approximately 4, 40, and 160 minutes of scanning time. Three subjects were used for pretraining base models, and a fourth subject was held out for adaptation. For the new subject, we applied fine-tuning followed by voxelwise regression ensemble using the limited data. As a baseline, we compared against the ”NRF scratch” approach, where a new NRF is trained entirely from the same limited dataset without pretraining. Across all data conditions, fine-tuning + ensemble consistently outperformed NRF scratch, confirming that NRF’s anatomically grounded formulation enables efficient cross-subject transfer, reducing the need for extensive subject-specific data while maintaining high predictive fidelity. The qualitative comparison is shown in Table 2. Predcition accuracy comparison across different methods is shown in Figure 2b. Notably, in the very low-data regime, finetuning + ensemble achieved strong semantic-level decoding performance. This shows that the strategy not only improves voxelwise prediction but also preserves subject variability, enabling predicted responses that more faithfully capture the semantic content of visual stimuli.
@@
Table 2: New subject adaptation with limited data (20, 200, 800 images). Here, the result is shown for using the NRF pretrained on subjects 1, 2, and 5 and adapted to subject 7 using finetuning + ensemble. This consistently outperforms training from scratch. With only 200 images, NRF outperforms the model trained with the whole dataset.
@@ 5.3 Probing Anatomical Awareness
NRF achieves strong performance, particularly in low-data regimes, by leveraging two key properties of fMRI data: (i) the local spatial continuity of voxel responses within a subject, and (ii) the anatomical alignment across subjects. To directly test the contribution of these factors, we design controlled perturbation experiments that disrupt either spatial smoothness or cross-subject correspondence. If NRF’s gains indeed stem from anatomical awareness, performance should degrade when these structural assumptions are broken.
Local Smoothness.
@@
To test whether NRF’s data efficiency stems from exploiting spatial continuity, we disrupted the natural smoothness of fMRI data by shuffling coordinate–response pairings. Voxel responses were randomly reassigned to MNI coordinates, breaking correlations between neighboring voxels. We performed two variants of this perturbation: (i) global shuffling, randomizing pairings across the entire visual cortex, and (ii) ROI-wise shuffling, randomizing only within each ROI. Traditional voxelwise models should be unaffected, since they treat voxels independently. In contrast, NRF relies on coordinate conditioning, and as expected, its performance dropped sharply in low-data regimes, with global shuffling producing the largest drop. This confirms that NRF’s improvements are driven by its ability to leverage local smoothness in brain responses. Shown in Figure 4(a).
Cross-Subject Alignment.
@@ 5.4 Ablation study
Voxelwise ensemble
A key component for new subject adaptation is voxelwise regression ensemble, where each voxel is fit with a linear regression model to optimally combine predictions from multiple fine-tuned base models. This approach improves prediction accuracy while preserving subject-specific variability. Table 3 compares voxelwise regression against single fine-tuned base models and simple averaging. While simple averaging slightly boosts voxelwise prediction accuracy, it hinders subject variability and produces predicted fMRI signals with reduced semantic fidelity, leading to lower decoding performance. In contrast, voxelwise regression leverages complementary information across base models in a flexible, voxel-specific way, achieving both higher voxel-level accuracy and stronger semantic-level decoding results.
Additional ablation results are included in the appendix A.5.
@@
Table 3: Ablation on voxelwise regression ensemble. We report the result for adapting from subjects 1,2,5 to subject 7 with 200 images.
@@ 6 Conclusion
In this work, we introduced the Neural Response Function (NRF), an anatomically aware neural encoding model that represents fMRI activity as a continuous function over MNI coordinates. Unlike conventional voxelwise models, NRF leverages spatial smoothness and cross-subject alignment to achieve accurate predictions in low-data regimes and to support efficient subject adaptation. Crucially, its continuous formulation moves beyond grid-locked voxels, allowing predictions at arbitrary spatial resolutions and across individuals. In this sense, NRF serves as a resolution-agnostic digital twin of the brain: a unified, flexible representation that integrates data across scales and subjects. These advances offer a new path toward efficient, generalizable, and anatomically grounded neural encoding.
@@ Appendix A Appendix
A.1 Use of LLMs
LLMs were used only for polishing grammar and writing clarity.
A.2 Additional details on NRF
@@
The predictor is a coordinate-conditioned MLP that takes as input both the Fourier positional encoding of the MNI coordinate and the fused image embedding $G(M)$. It outputs a single scalar—the predicted fMRI response at that voxel location. We use an 8-layer MLP with hidden dimension 4096, applying ReLU activations after each layer except the final output. Further ablations on model architecture are included in the appendix A.5.
A.3 Additional details on baseline models
@@
For linear regression baseline we used the same encoding model as Luo et al. (2023). Specifically, we extract the $(1\times 512)$ CLIP embedding from OpenAI CLIP ViT-B/16 and directly map it to the voxel dimension (e.g., 15,724 voxels) using a linear layer. The model is trained for 150 epochs with the AdamW optimizer, with a learning rate that decays linearly from $3\times 10^{-4}$. During inference, we select the checkpoint that achieves the lowest validation MSE.
A.4 Additional Details on evaluation metrics
@@
We used the evaluation metrics for decoded image evaluation from MindEye2 Scotti et al. (2024) directly. PixCorr measures the pixel-wise correlation between the ground-truth image and the reconstruction. SSIM refers to the Structural Similarity Index, which evaluates perceptual similarity between ground-truth and reconstructed images. Alex(2), Alex(5), Incep, and CLIP are two-way identification metrics (chance = 50%) based on feature similarity. Specifically, Alex(2) uses features from the 2nd layer of AlexNet, Alex(5) from the 5th layer of AlexNet, Incep from the final pooling layer of InceptionV3, and CLIP from the final layer of CLIP ViT-L/14. In two-way identification, the task is to decide whether the voxel embedding is closer to its paired image embedding or to a randomly selected image embedding, reported as percent correct. Eff and SwAV denote representational similarity metrics, computed as the average correlation distance between voxel embeddings and features extracted from EfficientNet-B1 and SwAV-ResNet50, respectively.
A.5 Additional ablation results
Finetuning strategy
Finetune strategy Voxel-Level Semantic-Level (via decoding)
@@ A.5 Additional ablation results
Both 0.234 0.361 0.254 0.377 0.910 0.958 90.9% 86.7% 0.730 0.424
Number of layers
Layers Voxel-Level Semantic-Level (via decoding)
@@
4 0.358 0.348 0.258 0.351 0.914 0.961 89.7% 85.1% 0.748 0.437
8 0.360 0.350 0.324 0.387 0.956 0.983 94.1% 89.9% 0.680 0.396
16 0.349 0.353 0.331 0.385 0.956 0.983 93.99% 89.9% 0.678 0.395
Hidden dimension
Hidden dimension Voxel-Level Semantic-Level (via decoding)

### v2

@@ 5 Results
Figure 2: Prediction accuracy (Pearson correlation $r$) in low data regime. a. Single-subject models. NRF consistently outperforms baseline models when trained on limited samples from scratch, highlighting the benefit of its continuous mapping.Results are shown for the mean of the median voxel correlation across four subjects, with error bars indicating the standard error of the mean (SEM). b. Cross-subject transfer. Voxel-level prediction accuracy visualized on the cortical surface of subject 7. When pretrained base models from other subjects are available, the NRF finetune ensemble further improves performance over NRF scratch and baselines, showing clear gains across visual regions.
@@ 5.1 Individual subjects encoding
We first evaluated NRF’s neural prediction capability for single-subject data. Training a separate model for each of the 4 subjects and comparing the average neural prediction accuracy across subjects. For comparison, we selected two representative encoding models as baseline comparison. The linear regression model from the BrainDIVE  (Luo et al., 2023) and the fWRF (Feature-Weighted Receptive Field) encoder  (St-Yves & Naselaris, 2018). Details about the baseline model are in Appendix A.2. We also took the result for full data encoding performance from MindSimulator  (Bao et al., 2025).
We first evaluate NRF under limited-data conditions, since practical applications rarely have access to the tens of thousands of trials collected in large-scale datasets such as NSD. As shown in Figure 2a, NRF achieves significantly higher accuracy than baseline models when trained on small numbers of images. Remarkably, with only 200 training samples, NRF outperforms baselines trained on more than 800 images. We attribute this data efficiency to the anatomical awareness of NRF: by conditioning on spatial coordinates, the model can exploit the smoothness of fMRI responses and learn more effectively from scarce data. This neuroscience-inspired design makes NRF particularly well-suited for realistic, low-data regimes.
Next, we evaluate NRF in the full-data setting($\sim$9k training images). Quantitative results, summarized in Table  1, show that NRF outperforms baselines on voxel-wise prediction metrics while achieving comparable performance on semantic-level evaluations. In addition, we observed that some baselines, such as fWRF, achieve unusually high semantic-level scores, in some cases even surpassing reconstructions from measured fMRI. We attribute this to decoder bias: fWRF outputs, while less neurally accurate, may align more closely with the pretrained decoder’s distribution, thereby inflating semantic metrics. These results indicate that semantic-level metrics should be interpreted as a coarse indication of reconstruction quality rather than a strict basis for comparing encoding models. Further discussion regarding these metrics is included in Appendix A.3.
Figure 3 provides a qualitative assessment via image reconstructions; the decoded images capture low-level visual features and high-level semantic categories with high fidelity to the ground truth. These results demonstrate that NRF maintains high voxel-level accuracy while also preserving semantic information, confirming its effectiveness across both limited and full data regimes. Appendix A.4 presents detailed per-subject results along with the voxel-wise accuracy distributions, and Appendix A.5 reports the statistical significance analyses.
@@
Table 1: Evaluation results of fMRI prediction accuracy for the model trained on the full dataset. All reported voxel-level metric values are reported as the per-subject median across voxels, and the table shows the mean of these medians across the 4 subjects. Additional subject-wise results and full per-voxel distributions are provided in Appendix A.4.
@@ 5.2 New Subject adaptation
More importantly, NRF enables cross-subject transfer, allowing knowledge learned from one subject to be adapted to new subjects—a critical property given that collecting fMRI data for new individuals is both resource-intensive and time-consuming. To evaluate this capability, we tested adaptation with 20, 200, and 800 images, corresponding to approximately 4, 40, and 160 minutes of scanning time. Three subjects were used for pretraining base models, and a fourth subject was held out for adaptation. For the new subject, we applied fine-tuning followed by voxel-wise regression ensemble using the limited data. As a baseline, we compared against the ”NRF scratch” approach, where a new NRF is trained entirely from the same limited dataset without pretraining. Across all data conditions, fine-tuning + ensemble consistently outperformed NRF scratch, confirming that NRF’s anatomically grounded formulation enables efficient cross-subject transfer, reducing the need for extensive subject-specific data while maintaining high predictive fidelity. The qualitative comparison is shown in Table 2. Prediction accuracy comparison across different methods is shown in Figure 2b. Notably, in the very low-data regime, finetuning + ensemble achieved strong semantic-level decoding performance. This shows that the strategy not only improves voxel-wise prediction but also preserves subject variability, enabling predicted responses that more faithfully capture the semantic content of visual stimuli.
@@
Table 2: New-subject adaptation with limited data (20, 200, 800 images). Results are shown for adapting the NRF pretrained on subjects 1, 2, and 5 to subject 7 using finetuning + ensemble. Performance is reported as the median across all voxels of subject 7 for voxel-level metrics. The adapted NRF consistently outperforms training from scratch, and with only 200 images, it even exceeds the performance of a model trained on the full dataset.
@@ 5.3 Probing Anatomical Awareness
To verify the mechanisms behind NRF’s performance, particularly in data-constrained settings, we probe its reliance on two fundamental properties of fMRI data: (i) the local spatial continuity of voxel responses within a subject, and (ii) the anatomical alignment across subjects. We conduct controlled ablation experiments by introducing structural perturbations that systematically disrupt these factors. If NRF’s efficiency stems from its anatomical grounding, performance should degrade significantly when these structural priors are violated.
Disrupting Local Smoothness via Voxel Shuffling.
@@
To test whether NRF’s data efficiency stems from exploiting spatial continuity, we disrupted the natural smoothness of fMRI data by shuffling coordinate–response pairings. Voxel responses were randomly reassigned to MNI coordinates, breaking correlations between neighboring voxels. We performed two variants of this perturbation: (i) global shuffling, randomizing pairings across the entire visual cortex, and (ii) ROI-wise shuffling, randomizing only within each ROI. Traditional voxel-wise models should be unaffected, since they treat voxels independently. In contrast, NRF relies on coordinate conditioning, and as expected, its performance dropped sharply in low-data regimes, with global shuffling producing the largest drop. This confirms that NRF’s improvements are driven by its ability to leverage local smoothness in brain responses. Shown in Figure 4(a).
Disrupting Cross-Subject Alignment.
@@ 5.4 Ablation study
voxel-wise ensemble
A key component for new subject adaptation is voxel-wise regression ensemble, where each voxel is fit with a linear regression model to optimally combine predictions from multiple fine-tuned base models. This approach improves prediction accuracy while preserving subject-specific variability. Table 3 compares voxel-wise regression against single fine-tuned base models and simple averaging. While simple averaging slightly boosts voxel-wise prediction accuracy, it hinders subject variability and produces predicted fMRI signals with reduced semantic fidelity, leading to lower decoding performance. In contrast, voxel-wise regression leverages complementary information across base models in a flexible, voxel-specific way, achieving both higher voxel-level accuracy and stronger semantic-level decoding results. This suggests that the ensemble does not merely act as a noise-reduction filter but effectively ”mixes” specialized knowledge from different source subjects to better represent the unique functional profile of the target individual. Additional ablation results are included in the Appendix A.7.
@@
Table 3: Ablation on voxel-wise regression ensemble. We report the result for adapting from subjects 1,2,5 to subject 7 with 200 images.
@@ 6 Conclusion
In this work, we introduced the Neural Response Function (NRF), an anatomically aware neural encoding model that represents fMRI activity as a continuous function over MNI coordinates. Unlike conventional voxel-wise models, NRF leverages spatial smoothness and cross-subject alignment to achieve accurate predictions in low-data regimes and to support efficient subject adaptation. Crucially, its continuous formulation moves beyond grid-locked voxels, allowing predictions at arbitrary spatial resolutions and across individuals. In this sense, NRF serves as a resolution-agnostic digital twin of the brain: a unified, flexible representation that integrates data across scales and subjects. These advances offer a new path toward efficient, generalizable, and anatomically grounded neural encoding.
@@ Appendix A Appendix
Sections
  1. 1.
    Additional details on NRF  A.1
  2. 2.
    Additional details on baseline models  A.2
  3. 3.
    Additional details on evaluation metrics  A.3
  4. 4.
    Additional results on single subject NRF  A.4
  5. 5.
    Statistical significance analysis  A.5
  6. 6.
    Additional results on new subject adaptation  A.6
  7. 7.
    Additional ablation results  A.7
  8. 8.
    High-Resolution NRF Querying A.8
A.1 Additional details on NRF
@@
The predictor is a coordinate-conditioned MLP that takes as input both the Fourier positional encoding of the MNI coordinate and the fused image embedding $G(M)$. It outputs a single scalar—the predicted fMRI response at that voxel location. We use an 8-layer MLP with hidden dimension 4096, applying ReLU activations after each layer except the final output. Further ablations on model architecture are included in Appendix A.7.
A.2 Additional details on baseline models
@@
For linear regression baseline we used the same encoding model as  (Luo et al., 2023). Specifically, we extract the $(1\times 512)$ CLIP embedding from OpenAI CLIP ViT-B/16 and directly map it to the voxel dimension (e.g., 15,724 voxels) using a linear layer. The model is trained for 150 epochs with the AdamW optimizer, with a learning rate that decays linearly from $3\times 10^{-4}$. During inference, we select the checkpoint that achieves the lowest validation MSE.
A.3 Additional Details on evaluation metrics
We used the evaluation metrics for decoded image evaluation from MindEye2  (Scotti et al., 2024) directly. Given an input image, we first predict its fMRI response using NRF; the predicted response is then fed into MindEye2, which reconstructs the corresponding visual stimulus. We compare these reconstructions against the ground-truth images presented during data collection.
@@ A.3 Additional Details on evaluation metrics
PixCorr measures the pixel-wise correlation between the ground-truth image and the reconstruction. SSIM refers to the Structural Similarity Index, which evaluates perceptual similarity between ground-truth and reconstructed images. Alex(2), Alex(5), Incep, and CLIP are two-way identification metrics (chance = 50%) based on feature similarity. Specifically, Alex(2) uses features from the 2nd layer of AlexNet, Alex(5) from the 5th layer of AlexNet, Incep from the final pooling layer of InceptionV3, and CLIP from the final layer of CLIP ViT-L/14. In two-way identification, the task is to decide whether the voxel embedding is closer to its paired image embedding or to a randomly selected image embedding, reported as percent correct. Eff and SwAV denote representational similarity metrics, computed as the average correlation distance between voxel embeddings and features extracted from EfficientNet-B1 and SwAV-ResNet50, respectively.
Interpretation of Semantic-Level Metrics
The semantic-level reconstruction metrics reported in the main paper (PixCorr, SSIM, AlexNet features, Inception, CLIP, EfficientNet, SwAV) are computed using the MindEye2 fMRI-to-image decoder, which reconstructs visual stimuli from predicted fMRI responses. We included these semantic metrics to provide additional qualitative insight into the behavior of the generated or selected images, following the evaluation format adopted in recent prior work, including MindSimulator (Bao et al., 2025), published at ICLR 2025. Our intention was to maintain consistency with existing evaluation practices, rather than to treat these metrics as decisive evidence for model comparison. While useful for assessing whether predicted responses support plausible reconstructions, these metrics therefore require careful interpretation.
Because MindEye2 is itself a trained neural decoder, it introduces its own representational biases. As a result, we observe counterintuitive cases in which fWRF predictions score higher than the actual measured fMRI responses on some semantic reconstruction metrics. This does not imply that fWRF predictions are more neuronally accurate; rather, it reflects that the decoder’s internal feature space is more compatible with certain statistical properties of the predicted responses than with the true fMRI signals.
This behavior demonstrates that semantic-level decoding metrics are not suitable as primary measures of encoding quality. Instead, they should be viewed as diagnostic tools for assessing whether predicted responses can support plausible image reconstructions.
Accordingly, all core model comparisons and statistical conclusions in the paper rely exclusively on voxel-level encoding metrics (e.g., Pearson correlation), which directly measure correspondence to ground-truth neural responses and are unaffected by decoder bias. Semantic reconstruction results are included for completeness and visualization, but should not be interpreted as primary indicators of model performance.
A.4 Additional results on single subject NRF
Here we report the single subject NRF evaluation for all subjects (Subject 1, 2, 5, 7) in table 4 and the Mean $\pm$ SEM across 4 subjects in table 5.
Voxel-Level Semantic-Level (via decoding)
Subject Method Pearson↑ MSE↓ PixCorr↑ SSIM↑ Alex(2)↑ Alex(5)↑ IncepT↑ CLIP↑ Eff↓ SwAV↓
Subject 1 Linear Regression 0.311 0.364 0.180 0.243 86.0% 94.3% 89.8% 84.2% 0.759 0.428
fWRF 0.341 0.363 0.304 0.342 96.8% 99.0% 96.2% 92.0% 0.615 0.357
NRF (ours) 0.361 0.349 0.324 0.387 95.6% 98.3% 94.1% 91.9% 0.680 0.396
Subject 2 Linear Regression 0.323 0.358 0.164 0.253 85.2% 94.2% 89.5% 83.9% 0.764 0.424
fWRF 0.347 0.364 0.219 0.222 96.9% 98.9% 95.3% 91.1% 0.635 0.359
NRF (ours) 0.368 0.347 0.240 0.351 89.3% 95.4% 88.1% 87.5% 0.767 0.442
Subject 5 Linear Regression 0.378 0.340 0.203 0.302 87.0% 97.6% 91.0% 85.2% 0.734 0.405
fWRF 0.413 0.357 0.345 0.403 96.7% 99.1% 96.9% 92.2% 0.599 0.352
NRF (ours) 0.425 0.346 0.236 0.379 93.5% 97.9% 97.1% 93.1% 0.609 0.317
Subject 7 Linear Regression 0.267 0.348 0.197 0.285 86.3% 94.1% 90.6% 84.7% 0.745 0.412
fWRF 0.269 0.360 0.346 0.399 97.4% 99.3% 96.3% 92.2% 0.609 0.357
NRF (ours) 0.278 0.328 0.244 0.367 88.0% 93.6% 89.2% 84.6% 0.768 0.445
Table 4: Single-subject NRF results for S1, S2, S5, and S7. For voxel-level metrics, the median values are reported.
Method Pearson↑ MSE↓ PixCorr↑ SSIM↑ Alex(2)↑ Alex(5)↑ IncepT↑ CLIP↑ Eff↓ SwAV↓